OpenStudy (anonymous):

1-2tan^2xcos^2=2cosx-1 prove te identity

OpenStudy (jdoe0001):

hmmm... is there a square in the right-hand side? \(\bf 1-tan^2(x)cos^2(x)=2cos^2(x)-1\quad ?\)

OpenStudy (jdoe0001):

hmm missing the 2... \(\bf 1-2tan^2(x)cos^2(x)=2cos^2(x)-1\quad ?\)

OpenStudy (anonymous):

yes there is 1-2tan^2xcos^2x=2cos^2x-1

OpenStudy (jdoe0001):

hmm do you recall what tangent = ?

OpenStudy (jdoe0001):

identity wise that is

OpenStudy (anonymous):

no

OpenStudy (jdoe0001):

http://www.mathwords.com/t/trig_identities.htm <--- can you see what tangent equals to?

OpenStudy (anonymous):

ok thank you

OpenStudy (jdoe0001):

\(\bf tan(x)=\cfrac{sin(x)}{cos(x)}\) thus \(\bf tan^2(x)=\cfrac{sin^2(x)}{cos^2(x)}\)

OpenStudy (anonymous):

right i get that part , but how do i incorporate the 2 in front of tan

OpenStudy (jdoe0001):

\(\bf 1-2tan^2(x)cos^2(x)\implies 1-2\left(\cfrac{sin^2(x)}{\cancel{cos^2(x)}}\right)\cancel{cos^2(x)}\\ \quad \\\implies 1-2sin^2(x)\)

OpenStudy (jdoe0001):

that's the left-side now let's take a peek at the right-side \(\bf \textit{recall that}\quad sin^2(x)+cos^2(x)=1\implies cos^2(x)=1-sin^2(x)\quad thus\\ \quad \\ 2cos^2(x)-1\implies 2[1-sin^2(x)]-1\)

OpenStudy (jdoe0001):

so.... what would you get for the right-side ?

OpenStudy (anonymous):

2cos^2(x)-1 right?

OpenStudy (jdoe0001):

\(\bf 2cos^2(x)-1\implies 2[1-sin^2(x)]-1\) multiply the 2 in front of the bracket, and cancel out like-terms

OpenStudy (anonymous):

ok i got it now thanks a lot you were extremely helpful

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

i have another question that is giving me difficulty